INFRARED TRIGGERED REACTION IN THE SF_{6}^{−}·HCOOH COMPLEX

\underline{HOLGER SCHNEIDER}, KAITO TAKAHASHI, REX T. SKODJE, J. MATHIAS WEBER, JILA, NIST, and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309.

Formic acid binds to the SF_{6}^{−} anion in a single (OH-F) H bond, with the CH group weakly tethered to a neighboring F atom. Similar to the case of SF_{6}^{−}· H_{2}O complexes, the SF bond involved in the (OH-F) H bond is significantly stretched and weakened by the attachment of the HCOOH ligand. The complex undergoes a reaction upon infrared absorption of one quantum in the OH or CH stretching mode of the formic acid moiety, leading predominantly to the formation of SF_{4}^{−}, HF, and CO_{2}. The reaction can be inhibited by attachment of two Ar atoms. We present IR photodissociation spectroscopy results and theoretical data illustrating the reaction mechanism.